OpenStudy (anonymous):

An arithmetic progression has first term a and common difference -1. The sum of the first n terms is equal to the sum of the first 3n terms. Express a in terms of n.

OpenStudy (jack1):

\[a _{n} = a \times k^{n-1}\] is that arithmetic progression formula?

OpenStudy (anonymous):

no its sn = n/2 (U1 +Un) or S(n)= n/2 (2u1 +(n-1)d)

OpenStudy (jack1):

no, ur right, mine was geometric, sorry arithmetic is \[a_n = a + d(n-1) \] a = first term d = common difference

OpenStudy (jack1):

so sum is as u described above @hartnn ...can u help with this plz?

hartnn (hartnn):

yes, so whats the sum of 1st n terms ? n/2 (2a -n+1) right ? what about sum of 3n terms ?

OpenStudy (jack1):

3n/2 (2a -3n+1)?

hartnn (hartnn):

yes, eddie agree ? you should also try, afterall its your question!

OpenStudy (anonymous):

yea i do

hartnn (hartnn):

good, so just equate them now n/2 (2a -n+1) = 3n/2 (2a -3n+1) notice that n/2 gets cancelled, after that its just a linear equation, can you simplify ?

OpenStudy (anonymous):

so I solve (2a -n+1) = 3 (2a -3n+1)

hartnn (hartnn):

yes, correct.

hartnn (hartnn):

but you got all the steps till here ?

OpenStudy (anonymous):

no I dont understand how you got there

hartnn (hartnn):

you have the sum of n terms formula, right ? just plug in d=-1, which is given, in that formula what u get ?

hartnn (hartnn):

and ofcourse, u1 =a

OpenStudy (anonymous):

I dont know what you're saying can you type it out

hartnn (hartnn):

S(n)= n/2 (2u1 +(n-1)d) is the sum formula put u1=a and d=-1 in this what u get ?

OpenStudy (anonymous):

S(n) = n/2( 2a + (n-1)-1) S(n) = n/2( 2a - n +1)

hartnn (hartnn):

yes, thats what i wrote for sum of n terms, now do the same thing for sum of 3n terms S(n)= n/2 (2u1 +(n-1)d) is the sum formula put u1=a and d=-1, n=3n in this what u get ?

OpenStudy (anonymous):

S(n) = 3n/2( 2a - 3n +1)

OpenStudy (anonymous):

ohhhhhhhhh

hartnn (hartnn):

yes, and since they are given to be equal, i just equated them.

OpenStudy (anonymous):

of so the final solution is 0= 2(2a-4n+1)

hartnn (hartnn):

how ? (2a -n+1) = 3 (2a -3n+1) (2a -n+1) = 6a -9n +3 continue... ?

OpenStudy (anonymous):

i did I made the problem equal to zero 2a -n+1 = 6a -9n +3 0= 4a -8n +2 0= 2(2a -4n +1)

hartnn (hartnn):

oh, okk..and u need to isolate a 2a =4n-1 a= ... ?

OpenStudy (anonymous):

a= 2n -1/2 !!!!!!!

hartnn (hartnn):

correct! :) good work.